Cub Company, a calendar-year entity, had 3,100 geothermal heating pumps in its beginning inventory for 20X1. On December 31, 20X0, the heating pumps had been adjusted down to $860 per unit from an actual cost of $980 per unit. It was the lower of cost or market. Cub purchased no additional units during 20X1. The following additional information is provided for 20X1:

Quarter Date Inventory
(units)
Unit Market
Value
1 March 31, 20X1 1,770 $ 845
2 June 30, 20X1 1,440 890
3 September 30, 20X1 1,320 800
4 December 31, 20X1 920 830


Required:
Respond to the following two independent scenarios as requested.

a. Case 1: The company does not have sufficient experience with the seasonal market for geothermal pumps and assumes that any reductions in market value during the year will be permanent.

(1) Determine the cost of goods sold for each quarter.a. Case 1: The company does not have sufficient experience with the seasonal market for geothermal pumps and assumes that anyb. Case 2: The company has prior experience with the seasonal market for geothermal pumps and expects that any reductions in
Answer #1

Answers Quarter 1, cost of goods solid = (1330 X860 ) + (15 X 1746) = 1143800 +26550 = 1170350 Quarter I cast of goods sold =cost of goods sold Quarta - $1120350 $241650 $ 182, 4.00 V Total $_1036400 $2660800 cost d pocks odd = ((3160–900) X 860) + (Quarter 3, cost of goods sold = 120 x $860 = $103200 Quartery cost of goods sold =(1330*4860) + (930x 30) = 1171400 Quarter |